A boat travels 7 km upstream and 7km back. The time for the round trip is 4 hrs. The speed of the stream is 3 ​km/hr. What is the speed of the boat in still​ water?
Answer by stanbon(75887)

A boat travels 7 km upstream and 7km back. The time for the round trip is 4 hrs. The speed of the stream is 3 ​km/hr. What is the speed of the boat in still​ water?
Upstream DATA:
dist = 7 km ; time = t hrs ; rate = 7/t km/hr
-----------
Downstream DATA:
dist = 7 km ; time = 4-t hrs ; rate = 7/(4-t) km/hr
-------
Rate Equations:
b - 3 = 7/t
b + 3 = 7/(4-t)
-------------------------
Solve for "t"::
[7/t]+3 = [7/(4-t)]-3
------
7/(4-t) - 7/t = 6
-------------------------
[7t-7(4-t)]/(4t-t^2) = 6
-------------------------------
14t-28 = 24t - 6t^2
--------------------------
6t^2 - 10t - 28 = 0
-------
3t^2 - 5t -14 = 0
t = 3.15 hrs
---
Solve for "b":
b-3 = 7/t
b = 7/3.15 + 3
boat speed = 5.22 km/hr
----------
Cheers,
Stan H.
---------------